OpenStudy (anonymous):

S(n) is a function that squares numbers. It can be written in equation form as S(n) = n2. What is the value of S(11)?

OpenStudy (anonymous):

What is \(11\) squared?

OpenStudy (anonymous):

s(11) = (11)^2

OpenStudy (anonymous):

You mean S(n)=n^2, if 11 = n, then it would be S(11) = 11^2